# SNN4Agents: A Framework for Developing Energy-Efficient Embodied Spiking Neural Networks for Autonomous Agents

### v1

@@ 6.4 Energy Efficiency Improvements
Experimental results for energy consumption are presented in Figure 14. These results show that, in general, our optimization techniques effectively reduce the energy consumption as compared to the baseline model (32b_20t). For instance, in the 100x100 attention window case, our optimizations lead to energy improvements by 1.36x - 1.41x for settings 16b_15t, 12b_15t, and 10b_15t; by 1.94x - 2.08x for settings 16b_10t, 12b_10t and 10b_10t; as well as by 3.82x - 4.03x for settings 16b_5t, 12b_5t, and 10b_5t; as shown by 9. Meanwhile, in the 50x50 attention window case, our optimizations lead to energy improvements by 1.35x - 1.54x for settings 16b_15t, 12b_15t, and 10b_15t; by 2.24x - 2.48x for settings 16b_10t, 12b_10t and 10b_10t; as well as by 4.32x - 4.66x for settings 16b_5t, 12b_5t, and 10b_5t; as shown by 10. These significant energy efficiency improvements come from the joint benefits from weight quantization, timestep reduction, and attention window reduction that decrease the processing power and processing latency, and hence the energy consumption.
Figure 14: Experimental results for energy consumption normalized to the baseline (32b_20t) across different parameter settings, including precision levels, timesteps, and training epochs under (a) 100x100 attention window, and (b) 50x50 attention window.
@@ 6.5 Trade-Off Analysis
Figure 15: Trade-off analysis for (a) accuracy and memory, (b) accuracy and normalized latency, and (c) accuracy and normalized energy consumption.
To properly select the appropriate SNN model for the given memory and latency constraints, we perform a trade-off analysis. To do this, we analyze the correlation between accuracy and memory footprint, accuracy and latency, as well as accuracy and energy consumption. For the accuracy-memory analysis, we observe that accuracy decreases as the memory footprint decreases, which is indicated by lower bit precision levels with smaller attention window sizes; see Figure 15(a). It means that we need to select the appropriate network model whose memory footprint meets the given memory constraint. For instance, if the memory constraint is 8MB, then we can select network models with 12- and 10-bit precision under 100x100 attention window, as well as all network models under 50x50 attention window. For the accuracy-latency analysis, we observe that accuracy decreases as the latency decreases, which is indicated by shorter timesteps; see Figure 15(b). It means that we need to select the appropriate network model whose latency meets the given memory constraint. For instance, if the latency constraint is 0.25x from the baseline latency, then we can select network models with 5 timestep under both 100x100 and 50x50 attention window sizes. For the accuracy-energy analysis, we observe that accuracy also decreases as the energy consumption decreases, which is indicated by lower bit precision levels with shorter timesteps and smaller attention window sizes; see Figure 15(c). For instance, if the given constraints are 8MB memory with 0.25x latency from the baseline latency, then we can select the network model with 10b_5t_100w setting, which achieves 84.12% with 68.75% memory saving and 3.58x speed-up, and 4.03x energy efficiency improvement from the baseline model; see 11. Furthermore, if the memory constraint is significantly smaller (e.g., 1MB memory) with 0.25x latency from the baseline latency, then we can select the network model with 10b_5t_50w setting, which achieves 77.10% with 97.32% memory saving and 4x speed-up, and 5.85x energy efficiency improvement from the baseline model; 12.
@@ 7 Conclusion
In this work, we propose an SNN4Agents framework that employs a set of optimization techniques for developing energy-efficient SNNs targeting autonomous agent applications. Here, our SNN4Agents compresses the SNN model through weight quantization, optimizes processing time through timestep reduction, and optimizes input samples through attention window reduction. The experimental results show that our proposed framework effectively improves energy efficiency, reduces memory footprint and latency, while maintaining high accuracy. If we consider a tolerance range of 2% lower accuracy from the baseline, we can achieve 84.12% accuracy with 68.75% memory saving, 3.58x speed-up, and 4.03x energy efficiency improvement. In this manner, our SNN4Agents enables the efficient development of SNN-based autonomous agents.
@@
Publicly available datasets were used in this study. These datasets can be found at the following site: https://www.prophesee.ai/2018/03/13/dataset-n-cars/ (NCARS dataset).
@@ Acknowledgments
This work has been supported in parts by the NYUAD Center for Cyber Security (CCS), funded by Tamkeen under the NYUAD Research Institute Award G1104, and Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.

### v2

@@ 6.4 Energy Efficiency Improvements
Experimental results for energy consumption are presented in Figure 14. These results show that, in general, our optimization techniques effectively reduce the energy consumption as compared to the baseline model (32b_20t). For instance, in the 100$\times$100 attention window case, our optimizations lead to energy improvements by 1.36x - 1.41x for settings 16b_15t, 12b_15t, and 10b_15t; by 1.94x - 2.08x for settings 16b_10t, 12b_10t and 10b_10t; as well as by 3.82x - 4.03x for settings 16b_5t, 12b_5t, and 10b_5t; as shown by 9. Meanwhile, in the 50$\times$50 attention window case, our optimizations lead to energy improvements by 1.35x - 1.54x for settings 16b_15t, 12b_15t, and 10b_15t; by 2.24x - 2.48x for settings 16b_10t, 12b_10t and 10b_10t; as well as by 4.32x - 4.66x for settings 16b_5t, 12b_5t, and 10b_5t; as shown by 10. These significant energy efficiency improvements come from the joint benefits from weight quantization, timestep reduction, and attention window reduction that decrease the processing power and processing latency, and hence the energy consumption.
Figure 14: Experimental results for energy consumption normalized to the baseline (32b_20t) across different parameter settings, including precision levels, timesteps, and training epochs under (a) 100$\times$100 attention window, and (b) 50$\times$50 attention window.
@@ 6.5 Trade-Off Analysis
Figure 15: (a) Trade-off analysis for accuracy vs. memory, accuracy vs. normalized latency, and accuracy vs. normalized energy consumption. (b) Computational complexity of different designs across different attention window sizes (i.e., 100$\times$100 and 50$\times$50) and timestep settings (i.e., 20, 15, 10, and 5).
To properly select the appropriate SNN model for the given memory and latency constraints, we perform a trade-off analysis. To do this, we analyze the correlation between accuracy and memory footprint, accuracy and latency, as well as accuracy and energy consumption. For the accuracy-memory analysis, we observe that accuracy decreases as the memory footprint decreases, which is indicated by lower bit precision levels with smaller attention window sizes; see Figure 15(a.1). It means that we need to select the appropriate network model whose memory footprint meets the given memory constraint. For instance, if the memory constraint is 8MB, then we can select network models with 12- and 10-bit precision under 100$\times$100 attention window, as well as all network models under 50$\times$50 attention window. For the accuracy-latency analysis, we observe that accuracy decreases as the latency decreases, which is indicated by shorter timesteps; see Figure 15(a.2). It means that we need to select the appropriate network model whose latency meets the given memory constraint. For instance, if the latency constraint is 0.25x from the baseline latency, then we can select network models with 5 timestep under both 100$\times$100 and 50$\times$50 attention window sizes. For the accuracy-energy analysis, we observe that accuracy also decreases as the energy consumption decreases, which is indicated by lower bit precision levels with shorter timesteps and smaller attention window sizes; see Figure 15(a.3). For instance, if the given constraints are 8MB memory with 0.25x latency from the baseline latency, then we can select the network model with 10b_5t_100w setting, which achieves 84.12% with 68.75% memory saving and 3.58x speed-up, and 4.03x energy efficiency improvement from the baseline model; see label- 11. Furthermore, if the memory constraint is significantly smaller (e.g., 1MB memory) with 0.25x latency from the baseline latency, then we can select the network model with 10b_5t_50w setting, which achieves 77.10% with 97.32% memory saving and 4x speed-up, and 5.85x energy efficiency improvement from the baseline model; see label- 12.
@@
6.6 Computational Complexity
Besides performance and efficiency benefits, a set of optimization techniques in our SNN4Agents framework also leads to different levels of computational complexity, which can be quantified through the number of synaptic and neuronal operations across different layers and the given timestep setting. Figure 15(b) shows the comparison of computational complexity from different designs across different attention window sizes (i.e., 100$\times$100 and 50$\times$50) and timestep settings (i.e., 20, 15, 10, and 5). In general, a smaller attention window incurs lower computational complexity than the bigger ones. The reason is that, a smaller attention window leads to smaller model size as indicated in Table 2 and Table 3, and hence smaller number of synaptic and neuronal operations. Specifically, if the larger attention window size is $W_{0}\times W_{0}$ and the smaller attention window size is $W_{1}\times W_{1}$, then the reduction factor of the computational complexity achieved by employing the smaller attention window can be estimated with $\approx(W_{0}/W_{1})^{2}$. For instance, the 20t_50w model has around 80% lower computation than the state-of-the-art (i.e., 20t_100w). Furthermore, a smaller number of timestep setting also incurs lower computational complexity than the bigger ones, as the synaptic and neuronal operations are performed in a timestep unit. Specifically, if the larger timestep is $T_{0}$ and the smaller timestep is $T_{1}$, then the reduction factor of the computational complexity achieved by employing the smaller timestep can be computed with $(T_{0}/T_{1})$. For instance, the 15t_100w, 10t_100w, and 5t_100w models incur lower computation than the 20t_100w by 25%, 50%, and 75%, respectively. Consequently, jointly employing smaller attention window size and smaller number of timestep will further decrease the computational complexity, whose reduction factor can be estimated with $\approx(W_{0}/W_{1})^{2}\cdot(T_{0}/T_{1})$. For instance, the 15t_50w, 10t_50w, and 5t_50w models incur lower computation than the 20t_100w by about 85%, 90%, and 95%, respectively.
6.7 Further Discussion
Our SNN4Agents framework is the first work that incorporates a set of optimization techniques while considering the event-based automotive dataset for enabling the efficient development of SNN-based autonomous agents. Therefore, it has several advantages as the following.
  • Our framework incorporates each optimization technique in a modular form. Therefore, the existing optimization techniques in the framework can be activated or deactivated as per the design requirements. Furthermore, new optimization techniques can also be incorporated feasibly into the framework.
  • Our framework can be used for design space exploration to investigate and understand the role of different SNN parameters. For instance, this framework can be utilized to observe the impact of specific parameters (e.g., membrane threshold potential) on the accuracy (Bano et al., 2024).
  • The generated SNN model has direct interfacing with event-based input data stream, thereby enabling efficient integration with event-based sensors (e.g., DVS Camera).
Besides these advantages, our framework in the current form can still be improved further to enhance the generated SNN model. Our SNN4Agents currently supports the STBP learning rule with offline-based training scenarios, which may be insufficient for some application use-cases. For instance, some autonomous agents may need to continuously adapt to dynamic operational environments, hence our SNN4Agents framework needs to be enhanced with advanced neural architectures and/or continual learning algorithms, while considering both offline and online training scenarios (Putra et al., 2024). Toward this, in the future, we plan to continue extending the work by incorporating more complex datasets in our framework, such as EventKITTI (Liang et al., 2022), then evaluating the performance as well as testing it for a real-world robotic application use-case (e.g., UGV rover).
@@ 7 Conclusion
In this work, we propose an SNN4Agents framework that employs a set of optimization techniques for developing energy-efficient SNNs targeting autonomous agent applications. Here, our SNN4Agents compresses the SNN model through weight quantization, optimizes processing time through timestep reduction, and optimizes input samples through attention window reduction. The experimental results show that our proposed framework effectively improves energy efficiency, reduces memory footprint and latency, while maintaining high accuracy. If we consider a tolerance range of 2% lower accuracy from the baseline, we can achieve 84.12% accuracy with 68.75% memory saving, 3.58x speed-up, and 4.03x energy efficiency improvement. In this manner, our SNN4Agents framework paves the way for further research and studies toward enabling the efficient development of SNN-based autonomous agents, and can be enhanced by incorporating other optimization techniques.
@@
Publicly available NCARS dataset can be accessed at https://www.prophesee.ai/2018/03/13/dataset-n-cars/. Meanwhile, codes of SNN4Agents are available at https://github.com/rachmadvwp/SNN4Agents.
@@ Acknowledgments
This work was partially supported by the NYUAD Center for Artificial Intelligence and Robotics (CAIR), funded by Tamkeen under the NYUAD Research Institute Award CG010.